A point charge of 3.0 nC with a mass of 4.0 g is moved from x=1.0m to x= 1.5m in an electric field of 5,0 N/C with the same direction as the motion of the charge. (a) How much work is done on the charge by the electric force? (b) What is the change in the potential energy of the charge? (c) Assuming that the charge started from rest, what is its speed at x=1.5 m?

 



1
Expert's answer
2022-02-20T08:20:35-0500

Given:

"q=3.0\:{\rm nC},\quad m=4.0\:\rm g"

"x_1=1.0\:{\rm m},\quad x_2=1.5\:{\rm m}"

"E=5.0\:\rm N/C"


(a)

"W=qEd\\
=3.0*10^{-9}*5.0*(1.5-1.0)=7.5*10^{-9}\:\rm J"

(b)

"V=\frac{W}{q}=\frac{7.5*10^{-9}}{3.0*10^{-9}}=2.5\:\rm V"

(c)

"v=\sqrt{2W/m}\\
=\sqrt{2*7.5*10^{-9}/4.0*10^{-3}}=1.9*10^{-3}\:\rm m/s"